OpenStudy (anonymous):

Find vector and parametric equations of this moving object: Velocity=(3,-1) and position at time t=0 is (2,3). Someone with vector knowledge! Please!

OpenStudy (amistre64):



OpenStudy (anonymous):

do you know how to find the vector and parametric equations?

OpenStudy (amistre64):

position is integration of velocity

OpenStudy (anonymous):

i'm sorry... i'm not sure what that means...

OpenStudy (amistre64):

that means im working on trying to understand the quesiton as posted :)

OpenStudy (amistre64):

is this velocity defined for all positions? or just the position of (2,3) ?

OpenStudy (anonymous):

Just that position, i assume. there's no other info given.

OpenStudy (amistre64):

if its defined for all positions; then we have a constant velocity to which every point is defined as: \[\int Velocity=\int(3,-1)=(3x+c_x,-y+c_y)\]

OpenStudy (amistre64):

or rather ts then xs and ys; kinda muddled a few things in the mind there lol

OpenStudy (amistre64):

but when t=0 then we know our constants .... they are 2 and 3

OpenStudy (amistre64):

so, r = <3t+2, -t+3> and this we would be able to work with forming equations

OpenStudy (mertsj):

www.mecca.org/~halfacre/MATH/plesson6.htm This could be exactly what you are looking for.

OpenStudy (amistre64):

this "freezing the system when you hit post" is getting annoying ....

OpenStudy (amistre64):

i would the assume that the parametrics are: x = 2 + 3t y = 3 - t

OpenStudy (mertsj):

That's what I think also.

OpenStudy (amistre64):

i never really use a "vector" form tho; it just seems pointless most of the time :)

OpenStudy (mertsj):

And then I think the vector equation is (2,3)+(3,-1)t

OpenStudy (amistre64):

it is, but some books like to use the <..> to encase the vector itself

OpenStudy (mertsj):

Yes. You are correct. I just didn't bother.

OpenStudy (mertsj):

I think you and I are more interested in this question than the asker.

OpenStudy (amistre64):

lol, at times we have to amuse ourselves ;)

OpenStudy (mertsj):

I guess.

OpenStudy (anonymous):

Lol i am pretty confused, but I think you explained it pretty well.... anyways. :??